OpenStudy (anonymous):

Which of the following is an equation of the line that passes through the point (950, 1,250) and has a slope of –10? A. y = –10x + 1,250 B. y = –10x + 10,750 C. y = –10x – 10,750 D. y = –10x – 8,250

OpenStudy (anonymous):

b

OpenStudy (harsimran_hs4):

y-1250 = -10(x - 950) y = -10x +9500 + 1250 y = -10x + 10750

OpenStudy (anonymous):

so i was right

OpenStudy (anonymous):

thank you

OpenStudy (harsimran_hs4):

Welcome!!